OpenStudy (anonymous):

Find the equation of the circle with center (2, -6) and radius 4

OpenStudy (anonymous):

equation of circle (x- h)^2 + (y - k)^2= r^2 where (h,k) is center & r is radius

OpenStudy (anonymous):

We have center (2,-6) & radius r=4 so just plug values in above equation ..

OpenStudy (anonymous):

Do u get it ?

OpenStudy (anonymous):

no it wasnt right but thanks for trying

OpenStudy (anonymous):

Actually, tgawade is correct. However, if you are looking for the general form of this equation, you will have to multiply it out. \[(x-2)^{2} + (y+6)^{2}=4^{2}\] The general form should look like this\[x ^{2} + y ^{2} -4x +12y +24= 0\]